# Interpretable Scaling Behavior in Sparse Subnetwork Representations of Quantum States

### v1

@@ 1 Introduction
Over the past decades, pruning has emerged as popular approach to reduce the number of parameters of neural networks lecun1989optimal ; hassibi1992second ; reed1993pruning ; han2015learning . The goal of pruning is to speed up model inference and improve portability without compromising on accuracy. In practice, contemporary models can be pruned to $50\%$ or even $90\%$ sparsity while mostly retaining their original performance gordon2020compressing ; gan2022playing ; yeo2023can . Pruning has also led to crucial insight on the inner workings of neural networks with many parameters: The lottery ticket hypothesis (LTH) states that sparse sub-networks (“winning tickets”) are contained within densely parametrized models that can match or exceed the performance of the original model when trained in isolation with proper initialization frankle2018lottery . The LTH has been verified in subsequent experimental and theoretical studies frankle2019stabilizing ; diffenderfer2021multi ; savarese2020winning ; chen2020long as well as extended to different types of neural networks yu2019playing ; chen2020lottery ; chen2021unified ; zhang2024graph .
We here examine the LTH in a non-traditional use-setting: The task of approximating the ground state of a quantum many-body system with a neural network carleo2017solving ; carrasquilla2017machine . Our task can be abstracted to the problem of learning a mapping from discrete data (spin-configurations $(\sigma_{1},...\sigma_{N})$, where $\sigma_{i}\in\{+1,-1\}$ and $N$ is the number of spin degrees of freedom) to a set of real numbers, often termed amplitudes, which specifies the ground state eigenvector of a many-body Hamiltonian.
The object we represent is profoundly different from that of typical machine learning tasks such as image recognition krizhevsky2017imagenet , or language modeling naveed2023comprehensive . There, a neural network is asked to represent extremely sparse objects, such as the subset of all RGB images that contain a cat, or the subset of random vocabulary strings that form coherent sentences. In contrast, a quantum many-body wavefunction is in general a fully dense representation, with probability amplitudes for different configurations spanning many orders of magnitudes. In order to capture the correct physics, these amplitudes must be encoded with high precision which typically necessitates the use of a more performant training algorithm than stochastic gradient descent. A common choice is natural gradient descent, which accounts for the curvature of the neural-network function.
@@
In order to look for robust principles, we vary both the neural network architecture and the physical Hamiltonian. Throughout this work, we consider two distinct physical models: the transverse field Ising model for different values of transverse field and the topological toric code model kitaev2003fault . We test the LTH on these cases and analyze the behavior of the pruned networks in terms of the physics of the ground state wavefunction.
Surprisingly, we find that an even stronger version of the LTH holds for the ground state search problem. By using a learned sparse structure with $5-20\%$ weights of the dense model, we are able to match the performance of a fully dense model using a random initialization of weights. We find that the optimal sparsity depends on the physics of the considered quantum many-body model. Additionally, we see that starting from the weights from the “lottery ticket initialization” offers negligible benefit over random weights. Crucially, the connectivity of the sparse masks by itself constitutes a valid lottery ticket independent of its initialization (see Section ( 4.1)).
In Section (4.2), we proceed to examine the scaling behavior of shallow FFNNs as a function of pruning and identify scaling regions that are universal with respect to varying initial network size, in accordance with pmlr-v139-rosenfeld21a : At both low and high sparsities, we find plateaus where the accuracy is fairly independent of the number of remaining parameters. Given that these plateaus occur at very different variational energies, they are likely representing different phases of matter. We gain further insights into the nature of these plateaus, and what happens in between, using tools to characterize phases of quantum matter. In particular, we show in Section (4.3) that the low-and high-error plateau correspond to physical phases of matter, connected by at least one (first-order) phase transition and possibly a third intermediate phase.
@@
  • We show that the LTH holds in a stronger version for neural-network representations of quantum many-body ground states: While we can prune up to $80-95\%$ sparsity without significant loss in accuracy, the performance of the so-obtained sparse subnetworks is largely independent of the parameter initialization when they are trained in isolation.
  • We demonstrate that neural network wavefunctions have distinct scaling regions across several quantum many-problems, consisting of high-and low error plateaus that bookend an approximately power law region. Overparameterization only affects the performance of the models in the low-error plateau.
@@ 2 Related work
Lottery ticket hypothesis and neural-network quantum states
In the context of neural-network quantum states, the effect of pruning has been studied for the task of reconstructing a known quantum many-body wave-function from data golubeva2022pruning ; sehayek2019learnability but the LTH itself has not been tested. Our work further differs by consideration of the distinct task of variational energy minimization for ground-state approximation and careful study of the pruning dynamics.
@@
This work is inspired by previous studies on predictable scaling laws bahri2024explaining of sparse neural networks during iterative pruning pmlr-v139-rosenfeld21a . We demonstrate the applicability of these methods by considering a non-traditional task of learning a quantum ground state. In addition the scaling behavior of dense neural network representations of quantum ground states chen2024empowering ; passetti2023can ; denis2023comment has been studied, but not in the context of pruning or other network-compression methods.
@@
Following the LTH, a stronger hypothesis ramanujan2020s was proven malach2020proving ; pensia2020optimal ; orseau2020logarithmic ; burkholz2021existence ; cunha2022proving ; ferbach2022general , showing that high-performing sparse subnetworks exist at initialization already for a variety of neural network architectures. The existence of strong lottery tickets suggest that the connectivity of these subnetworks may be highly dependent on the initial parameterization. Our evidence points to the contrary in the case of neural-network quantum states, i.e. independence with respect to parameter initialization. Other related studies have shown that parameter sign configurations play an important role in obtaining winning ticketsgadhikar2024masks ; zhou2019deconstructing ; oh2025find . We extend on these results by relating the sign structure to the ground state physics of the considered model.
@@
Several works have used tools from statistical mechanics in relation to neural network representations. One approach treats the neural network as a Hamiltonian and the weights as couplings barney2024neural ; winer2025deep . Another perspective focuses on phase transitions in the context of understanding a neural network’s ability to generalize as the number of training samples is increased sompolinsky1990learning ; gyorgyi1990first ; barbier2019optimal ; schwarze1993learning ; maillard2020phase ; barra2017phase ; cui2024phase . Here, we are able to quantify phases of a neural network in a very direct fashion: Our neural network wavefunction corresponds to a real quantum phase with measurable observables.
@@ 3.1 Neural network representations of quantum states
Throughout this work, we consider systems with $N$ spin-$1/2$ degrees of freedom. We aim to solve the ground-state problem. In particular, we are given a Hamiltonian $\hat{H}$ that describes the considered system and acts as operator on quantum states. The ground state $|\Psi\rangle_{G}$ corresponds to the lowest-lying eigenstate of the Hamiltonian, i.e.
@@
where $E_{G}$ is the system’s ground state energy. In general, the dimensionality of this problem renders an exact solution unfeasible: The Hilbert space size scales as $2^{N}$.
@@
Here, the (complex) wave-function amplitudes $\Psi^{NN}_{\theta}(\bm{\sigma})$ are parametrized with the help of a neural network: The output value of the network is chosen to represent $\log(\Psi_{\theta}^{NN}(\bm{\sigma}))$ when given as input a spin-configuration $\bm{\sigma}=(\sigma_{1},...\sigma_{N})$, $\sigma_{i}\in\{\pm 1\}$. The weights and biases of the network form the set of parameters $\theta$. Throughout this work, we only consider wave functions that are known to be real and positive. Our experiments are performed on three different network architectures: shallow feed-forward neural networks (FFNNs), shallow convolutional neural networks (CNNs) and few-layer residual convolutional neural networks (ResCNNs). More details on the employed neural networks and hyperparameters can be found in the Appendix.
@@
Here, $\langle\cdot\rangle_{\theta}$ denotes the expectation value taken with respect to the wave-function $|\Psi^{NN}_{\theta}\rangle$. All expectation values are computed via Monte Carlo sampling. We minimize this cost function via natural gradient descent, in the context of the ground-state search also termed “stochastic reconfiguration” sorella1998green . For more details on sampling and optimization, see Appendix. The simulations are carried out with the help of the package netket carleo2019netket ; vicentini2022netket , based on JAX bradbury2018jax and FLAX flax2020github .
@@
where the reference exact ground-state energy $E_{\mathrm{exact}}$ is computed either via exact diagonalization on a small system, the density-matrix-renormalization group (DMRG) or the lowest energy that was obtained with our neural network quantum state. In some cases throughout this work, we choose to report the absolute error per spin $\Delta E=|E_{\mathrm{exact}}-\langle\hat{H}\rangle_{\theta}|/N$, when comparing errors in different quantum phases or across multiple system sizes. All DMRG computations are carried out using the package ITensor itensor . The implementation details can be found in the Appendix.
@@ 3.2 Physical models
with the Pauli matrices $\hat{\sigma}^{x/z}$. The interplay between spin coupling of nearest neighboring spins (denoted by $\langle i,j\rangle$) and the transverse magnetic field $\kappa>0$ determines the quantum phase of the model. In particular, the Ising model exhibits a quantum phase transition as a function of magnetic field $\kappa$: For small magnetic field, the spin coupling term dominates and the ground state is ferromagnetic. In the case of a strong transverse field $\kappa$, spins are aligned in form of a paramagnet in the $+x$ direction. The quantum phase transition is estimated to occur at $\kappa_{c}\approx 3.04438(2)$ PhysRevE.66.066110 . At the position of the phase transition, the system is at criticality and spin-spin correlation functions decay with power-law.
@@
As a use-case of a model with an exact solution and topological order, we consider the toric code model kitaev2003fault . The toric code model is defined on a $L\times L$ lattice with periodic boundary conditions and $N=2L^{2}$ spin-$1/2$ degrees of freedom on the edges. The Hamiltonian is given by
@@
where $\hat{A}_{v}=\prod_{i\in v}\hat{\sigma}^{x}_{i}$ and $\hat{B}_{p}=\prod_{i\in p}\hat{\sigma}^{z}_{i}$ act on the plaquettes and vertices on the given lattice. The ground-state manifold is four-fold degenerate and consists of eigenstates of the mutually commuting $\hat{A}_{v}$ and $\hat{B}_{p}$ with eigenvalue $+1$. The toric code serves as instructive example to test interpretability of neural-network quantum states, as several exact constructions of the toric code ground states with the help of neural networks have been brought forward in previous work deng2017machine ; carrasquilla2017machine ; valenti2022correlation ; chen2025representing .
@@ 3.3 Sparse training strategies
Concretely, the IMP-WR procedure begins with training a randomly initialized dense neural network (pre-training phase). The optimized parameters $\theta_{\mathrm{wr}}$ from this pre-training phase serve as rewinding point: Once this rewinding point is obtained, the IP-WR procedure proceeds in an iterative fashion. In each pruning iteration, the network’s remaining non-zero weights are initialized in the rewinding point and then trained. After training, a fraction $p$ of the weights with smallest magnitude removed in an unstructured fashion. We use a pruning ratio of $p_{r}=0.12$ for FFNNs and ResCNNs, and a smaller pruning ratio of $p_{r}=0.05$ for CNNs due to their initially lower parameter count. This process is repeated for a total of $I$ pruning iterations. The number of remaining parameters $n$ of the neural network at a particular iteration $i$ is given by
@@
The $T({\theta}_{\mathrm{init}},m_{\mathrm{imp}})$ variant tests the original LTH, while $T({\theta}_{\mathrm{rand}},m_{\mathrm{imp}})$ tests the viability of using the same mask on a new parameter initialization. Finally, $T({\theta}_{\mathrm{init}},m_{\mathrm{rand}})$ is a random sparse initialization of the pruning mask, which serves as a benchmark to the other two isolated training variants.
@@ 4.1 Testing the lottery ticket hypothesis
The original LTH frankle2018lottery asserts that it is the combination of masks and initial parameter values that forms a winning ticket. Thus we would expect training $T(\theta_{\mathrm{init}},m_{\mathrm{imp}})$ in isolation to greatly outperform the other ticket initialization choices. While we find that winning ticket networks can replicate the performance of the dense network (up to a critical sparsity), we also observe that the same sparse masks perform equally well when parameters are randomly initialized. On the other hand, we find that sparse networks based on masks with randomly chosen connectivity perform significantly more poorly; the performance begins to degrade with a small amount of pruning, and they reach high error much faster (see Fig.  1).
Figure 1: Test of the LTH at the critical point $\kappa=\kappa_{c}$ of the TFIM on the $N=10\times 10$ lattice, for various neural network architectures: (a) shallow CNN, (b) shallow FFNN, and (d) Res-CNN (shaded regions are statistical sampling errors). The relative error is computed with respect to the overall lowest obtained energy, here the average energy of the three best-performing dense ResCNNs (by smallest variance) as shown in the inset of (b), which produce a smaller energy than DMRG with a bond dimension $\chi=3000$.

### v2

# Connectivity determines the capability of sparse neural network quantum states

@@ 1 Introduction
Over the past decades, pruning has emerged as popular approach to reduce the number of parameters of neural networks LeCun et al. (1989); Hassibi and Stork (1992); Reed (1993); Han et al. (2015). The goal of pruning is to speed up model inference and improve portability without compromising on accuracy. In practice, contemporary models can be pruned to $50\%$ or even $90\%$ sparsity while mostly retaining their original performance Gordon et al. (2020); Gan et al. (2022); Yeo et al. (2023). Pruning has also led to crucial insight on the inner workings of neural networks with many parameters: The lottery ticket hypothesis (LTH) states that sparse sub-networks (“winning tickets”) are contained within densely parametrized models that can match or exceed the performance of the original model when trained in isolation with proper initialization Frankle and Carbin (2018). The LTH has been verified in subsequent experimental and theoretical studies Frankle et al. (2019); Diffenderfer and Kailkhura (2021); Savarese et al. (2020); Chen et al. (2020b) as well as extended to different types of neural networks Yu et al. (2019); Chen et al. (2020a, 2021); Zhang et al. (2024).
We here examine the LTH in a non-traditional use-setting: The task of approximating the ground state of a quantum many-body system with a neural network Carleo and Troyer (2017); Carrasquilla and Melko (2017). Our task can be abstracted to the problem of learning a mapping from discrete data (spin-configurations $(\sigma_{1},...\sigma_{N})$, where $\sigma_{i}\in\{+1,-1\}$ and $N$ is the number of spin degrees of freedom) to a set of real numbers, called amplitudes, which specifies the ground state eigenvector of a many-body Hamiltonian.
The object we represent is profoundly different from that of typical machine learning tasks such as image recognition Krizhevsky et al. (2017), or language modeling Naveed et al. (2023). There, a neural network is asked to represent extremely sparse objects, such as the subset of all RGB images that contain a cat, or the subset of random vocabulary strings that form coherent sentences. In contrast, a quantum many-body wavefunction is in general a fully dense representation, with probability amplitudes for different configurations spanning many orders of magnitudes. In order to capture the correct physics, these amplitudes must be encoded with high precision which typically necessitates the use of a more performant training algorithm than stochastic gradient descent. A common choice is natural gradient descent, which accounts for the curvature of the neural-network function.
@@
In order to look for robust principles, we vary both the neural network architecture and the physical Hamiltonian. Throughout this work, we consider two distinct physical models: the transverse field Ising model for different values of transverse field and the topological toric code model Kitaev (2003). We test the LTH on these cases and analyze the behavior of the pruned networks in terms of the physics of the ground state wavefunction.
Surprisingly, we find that an even stronger version of the LTH holds for the ground state search problem on the Hamiltonians we consider in this work. By using a learned sparse structure with $5-20\%$ weights of the dense model, we are able to match the performance of a fully dense model using a random initialization of weights. We find that the optimal sparsity depends on the physics of the considered quantum many-body model. Additionally, we see that starting from the weights from the “lottery ticket initialization” offers negligible benefit over random weights. Crucially, the connectivity of the sparse masks by itself constitutes a valid lottery ticket independent of its initialization (see Section ( 4.1)).
In Section (4.2), we proceed to examine the scaling behavior of shallow FFNNs as a function of pruning and identify scaling regions that are universal with respect to varying initial network size, in accordance with Rosenfeld et al. (2021): At both low and high sparsities, we find plateaus where the accuracy is fairly independent of the number of remaining parameters. Given that these plateaus occur at very different variational energies, they are likely representing different phases of matter. We gain further insights into the nature of these plateaus, and what happens in between, using tools to characterize phases of quantum matter. In particular, we show in Section (4.3) that the low-and high-error plateau correspond to physical phases of matter, connected by at least one (first-order) phase transition and possibly a third intermediate phase.
@@
  • We show that neural network quantum many-body wavefunctions have a different form of the LTH. Here, the lottery tickets are sparse subnetworks, and the initial weights are not relevant to the final performance. This contrasts with the LTH in other problem domains Frankle and Carbin (2018); Chen et al. (2020a), where a lottery ticket consists of a sparse subnetwork initialized with particular parameter values.
  • We demonstrate that neural network wavefunctions have distinct scaling regions across several quantum many-problems, consisting of high-and low error plateaus that bookend an approximately power law region. Changing the number of parameters only affects the performance of the models in the low-error plateau.
@@ 2 Related work
Pruning neural-network quantum states
In the context of neural-network quantum states, the effect of pruning has been studied for the task of reconstructing a known quantum many-body wave-function from a 1D Hamiltonian Golubeva and Melko (2022); Sehayek et al. (2019), which is a supervised learning task. In contrast, we study an unsupervised learning task of finding the minimum energy variational wavefunction of 2D Hamiltonians. Our work also interprets the structure of the learned sparse representations in light of the physical Hamiltonian.
@@
This work is inspired by previous studies on predictable scaling laws Bahri et al. (2024) of sparse neural networks during iterative pruning Rosenfeld et al. (2021). We test these methods on a non-traditional task of learning a quantum ground state. For neural quantum states, it is important to understand how the number of parameters necessary to achieve a desired relative error scales with system size. While this relation has been studied for dense models Chen and Heyl (2024); Passetti et al. (2023); Denis et al. (2023), here we examine scaling behavior through the lens of sparse networks by pruning initially dense models and analyzing the resulting error scaling.
@@
Following the LTH, a stronger hypothesis Ramanujan et al. (2020) was proven Malach et al. (2020); Pensia et al. (2020); Orseau et al. (2020); Burkholz et al. (2021); da Cunha et al. (2022); Ferbach et al. (2022), showing that high-performing sparse subnetworks already exist at initialization (“strong lottery tickets") for a variety of neural network architectures. Traditionally, strong lottery tickets are therefore characterized by both a sparse subnetwork and a set of initial parameter values. In support of this, it has also been shown in the original LTH that changing the initialization of lottery tickets results in a performance decrease Frankle and Carbin (2018). Our evidence points to the contrary for the case of neural-network quantum states, as the lottery ticket performance only depends on its structure, not its parameter initialization. Other related studies have shown that parameter sign configurations play an important role in obtaining winning tickets Gadhikar and Burkholz (2024); Zhou et al. (2019); Oh et al. (2025). In this work, we are able to find sparse subnetwork representations that correspond to asymptotically exact solutions of the problem at hand in both absolute values of the weights as well as their sign structure.
@@
Several works have used tools from statistical mechanics in order to analyze neural network representations. One approach understands phase transitions as a function of training time, by considering the neural network as a classical Hamiltonian with individual neurons as spin variables, and the weights as time-dependent coupling strengths as a function of training under stochastic gradient descent Barney et al. (2024); Winer and Hanin (2025). Another perspective characterizes phase transitions in a neural network’s generalization capabilities as a function of training samples by employing replica methods from statistical mechanics Sompolinsky et al. (1990); Györgyi (1990); Barbier et al. (2019); Schwarze (1993); Maillard et al. (2020); Barra et al. (2017); Cui et al. (2024). Our work makes a direct connection to phase transitions in quantum many-body physics – we characterize phases directly by computing order parameters on a many-body wavefunction parameterized by a neural network. We find that these order parameters can change as a function of sparsity.
@@ 3.1 Neural network representations of quantum states
Throughout this work, we consider systems with NN spin-$1/2$ degrees of freedom. We aim to solve the ground-state problem. In particular, we are given a Hamiltonian $\hat{H}$ for which we want to find the lowest-lying eigenstate, i.e.
@@
where $E_{G}$ is the system’s ground state energy. In general, the dimensionality of this problem renders an exact solution unfeasible as the Hilbert space size scales as $2^{N}$.
@@
Here, the wave-function amplitudes $\Psi^{NN}_{\theta}(\bm{\sigma})$ are parametrized with the help of a neural network: The output value of the network is chosen to represent $\log(\Psi_{\theta}^{NN}(\bm{\sigma}))$ when given as input a spin-configuration $\bm{\sigma}=(\sigma_{1},...\sigma_{N})$, $\sigma_{i}\in\{\pm 1\}$. The weights and biases of the network form the set of parameters $\theta$. Throughout this work, we only consider wave functions that are known to be real and positive. Our experiments are performed on three different network architectures: shallow feed-forward neural networks (FFNNs), shallow convolutional neural networks (CNNs) and few-layer residual convolutional neural networks (ResCNNs). More details on the employed neural networks and hyperparameters can be found in the Appendix.
@@
Here, $\langle\cdot\rangle_{\theta}$ denotes the expectation value taken with respect to the wave-function $|\Psi^{NN}_{\theta}\rangle$. All expectation values are computed via Monte Carlo sampling. We minimize this cost function via natural gradient descent, in the context of the ground-state search also termed “stochastic reconfiguration” (SR) Sorella (1998). For more details on sampling and optimization, see Appendix. The simulations are carried out with the help of the package netket Carleo et al. (2019); Vicentini et al. (2022), based on JAX Bradbury et al. (2018) and FLAX Heek et al. (2024).
@@
where the reference exact ground-state energy $E_{\mathrm{exact}}$ is computed either via exact diagonalization on a small system, the density-matrix-renormalization group (DMRG) or the lowest energy that was obtained with our neural network quantum state. In some cases throughout this work, we choose to report the absolute error per spin $\Delta E=|E_{\mathrm{exact}}-\langle\hat{H}\rangle_{\theta}|/N$, when comparing errors in different quantum phases or across multiple system sizes. All DMRG computations are carried out using the package ITensor Fishman et al. (2022). The implementation details can be found in the Appendix.
@@ 3.2 Physical models
with the Pauli matrices $\hat{\sigma}^{x/z}$. The interplay between spin coupling of nearest neighboring spins (denoted by $\langle i,j\rangle$) and the transverse magnetic field $\kappa>0$ determines the quantum phase of the model. In particular, the Ising model exhibits a quantum phase transition as a function of magnetic field $\kappa$: For small magnetic field, the spin coupling term dominates and the ground state is ferromagnetic. In the case of a strong transverse field $\kappa$, spins are aligned in form of a paramagnet in the +x+x direction. The quantum phase transition is estimated to occur at $\kappa_{c}\approx 3.04438(2)$ Blöte and Deng (2002). At the position of the phase transition, the system is at criticality and spin-spin correlation functions decay with power-law.
@@
As a use-case of a model with an exact solution and topological order, we consider the toric code model Kitaev (2003). The toric code model is defined on a $L\times L$ lattice with periodic boundary conditions and $N=2L^{2}$ spin-$1/2$ degrees of freedom on the edges. The Hamiltonian is given by
@@
where $\hat{A}_{v}=\prod_{i\in v}\hat{\sigma}^{x}_{i}$ and $\hat{B}_{p}=\prod_{i\in p}\hat{\sigma}^{z}_{i}$ act on the plaquettes and vertices on the given lattice. The ground-state manifold is four-fold degenerate and consists of eigenstates of the mutually commuting $\hat{A}_{v}$ and $\hat{B}_{p}$ with eigenvalue +1+1. The toric code serves as instructive example to test interpretability of neural-network quantum states, as several exact constructions of the toric code ground states with the help of neural networks have been brought forward in previous work Deng et al. (2017); Carrasquilla and Melko (2017); Valenti et al. (2022); Chen et al. (2025b).
@@ 3.3 Sparse training strategies
Concretely, the IMP-WR procedure begins with training a randomly initialized dense neural network (pre-training phase). The optimized parameters $\theta_{\mathrm{wr}}$ from this pre-training phase serve as rewinding point: Once this rewinding point is obtained, the IP-WR procedure proceeds in an iterative fashion. In each pruning iteration, the network’s remaining non-zero weights are initialized in the rewinding point and then trained. After training, a fraction $p_{r}$ of the weights with smallest magnitude removed in an unstructured fashion. We use a pruning ratio of $p_{r}=0.12$ for FFNNs and ResCNNs, and a smaller pruning ratio of $p_{r}=0.05$ for CNNs due to their initially lower parameter count. This process is repeated for a total of $I$ pruning iterations. The number of remaining parameters $n$ of the neural network at a particular iteration $i$ is given by
@@
The $T({\theta}_{\mathrm{init}},m_{\mathrm{imp}})$ variant tests the original LTH, while $T({\theta}_{\mathrm{rand}},m_{\mathrm{imp}})$ tests the viability of using the same mask on a new parameter initialization. Finally, $T({\theta}_{\mathrm{init}},m_{\mathrm{rand}})$ is a random sparse initialization of the pruning mask, which serves as a benchmark to the other two isolated training variants. The isolated training variants are tested with masks derived from single trajectories of iterative magnitude pruning, and we report the statistics over several random initializations in the Appendix.
@@ 4.1 Testing the lottery ticket hypothesis
The original LTH Frankle and Carbin (2018) asserts that it is the combination of masks and initial parameter values that forms a winning ticket. Thus we would expect training $T(\theta_{\mathrm{init}},m_{\mathrm{imp}})$ in isolation to greatly outperform the other ticket initialization choices. While we find that winning ticket networks can replicate the performance of the dense network (up to a critical sparsity), we also observe that the same sparse masks perform equally well when parameters are randomly initialized. On the other hand, we find that sparse networks based on masks with randomly chosen connectivity perform significantly more poorly; the performance begins to degrade with a small amount of pruning, and they reach high error much faster (see Fig.  1).
Figure 1: Test of the LTH at the critical point $\kappa=\kappa_{c}$ of the TFIM on the $N=10\times 10$ lattice, for various neural network architectures: (a) shallow CNN, (b) shallow FFNN, and (c) Res-CNN (shaded regions are statistical sampling errors). The relative error is computed with respect to the average energy of three best-performing dense ResCNNs by smallest variance (see Appendix).